OpenStudy (anonymous):

I WILL GIVE MEDAL AND FOLLOW AND WRITE A TESTIMONIAL Is the simplified form of 2 to the square root of 3 times the square root of 12 rational?

OpenStudy (igreen):

\(2 \sqrt{3} \times \sqrt{12}\)

OpenStudy (anonymous):

Yes how did u do the square root sign lol

OpenStudy (igreen):

Click the equation button right below your chat box :)

OpenStudy (igreen):

And your answer is 12 :)

OpenStudy (anonymous):

Is it rational tho ?

OpenStudy (igreen):

You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

@iGreen is it rational tho

OpenStudy (igreen):

The rational answer is \(\sqrt{144}\), because the Square root of 144 is 12.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Thx im a fan i wrote testimonial and gave u a medak